A Bibliography of Publications in the IEEE symposia on Foundations of Computer Science (FOCS) for

1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 2.06

Title word cross-reference

(0, 1) [HLW88]. (t + 1) [MW88]. 0 [PS89]. 1 [PS89]. 1/2 + 1P oly(log N) [ACGS84]. 2 [DS89b]. g [Mal88]. K [Vai84, Fri84, GH88, PSS89]. log^{c} n [BR89b]. n [Fri84]. O(g) [Mal88]. O(n log n) [Fri84]. ω [Saf88]. s [KS89]. t [KS89].

* [BV89]. *-tree [BV89].

-automata [Saf88]. -cut [GH88]. -Dimensional [Vai84]. -functions [HLW88]. -round [MW88]. -sets [PSS89]. -th [Fri84]. -way [DS89b]. -wise [BR89b].

21st [IEE80]. 22nd [IEE81]. 23rd [IEE82]. 24th [IEE83]. 25th [IEE84, Ano84]. 26th [IEE85]. 27th [IEE86]. 28th [IEE87]. 29th [IEE88].

30th [IEE89].

abstract [Ong88]. Abstraction [Mul84, Rus89]. Achieving [CK88]. addresses [BAG88]. Agreement [BD84, MW88]. algebraic [Yao89]. Algorithm [Col84, FP84, KMR84, Mis84, SG84, Sul84, Vai84, CE88, Cha89, CH89, Cla88, DHK88, DK89, GH88, Jus89, LW89, Mul88, RR89, Ren88, Vai89a]. Algorithms [BCLS84, FT84, GGS84, GW84, IPK84, Sho84, AA88, BRS89, DH89, GX89, GPT88, GPV88, Han89, KS89, KMV89, Kle88, LR88, LMR88a, LM89, MS89, MNN89, Sho88]. alternating [LL89]. amplification [CW89]. analogue [FL89]. Analysis [FHV84, Sho84]. Angeles [IEE87]. Annual [Ano84, IEE81, IEE83, IEE80, IEE82, IEE84, IEE85, IEE86, IEE87, IEE88, IEE89]. Applications [IA84, KMS84, MSM84, BRS89, DS89a, GK89, Han89, LR88, Sie89]. Approach [IR84]. approximate [LR88]. approximating [MST89]. Approximation [HS89, Han89, Vai84, AA88, LR88]. arbitrary [DK89, Jus89]. Area [KWW89]. Area-optimal [KWW89]. Arithmetic [Ebe84, Mil89]. Arizona [IEE83]. arrangements [CEG^{+}88]. Arrays [FS84, AP88]. assignment [GX89]. associative [KP88]. assumptions [CK88]. Asymptotically [DH89]. asynchronous [CM89]. Augmenting [SG84]. automata [DS89b, EJ88, Jia89, Saf88, Tse89]. automorphism [Dic89]. Average [BS84b, BCLS84, Sho84, LV89a]. Average-Case [Sho84].

based [Hau89, IL89a]. basic [CK89]. basis [FL89]. be [LM89]. Behavior [BCLS84]. Biased [Blu84]. Biconnected [TV84]. Bidirectional [Mis84]. Bin [Sho84]. binary [GRS89]. Bisection [BCLS84]. Bits [ACGS84, BM82, Cai89, IZ89]. black [KT88]. Boolean [Fri84, KKL88, RW89]. Bound [GW84, Bsh88, CFI89, PSS89]. bounded [CL89b, Her89]. Bounds [BK88, Tiw84, AGR89, Cai89, CEG^{+}88, DKM^{+}88, Hau89, KGY89, MST88, Ng89, OY88, Yao89]. boxes [KT88]. branching [CL89a]. Buffers [Pip84]. Byzantine [BD84, MW88].

calculating [Mil89]. Calculi [Mul84]. Calculus [Mit84, Ong88]. California [IEE87]. can [Koc88]. Canada [IEE86].

Carolina [IEE89]. cascades [Sun89]. Case [BCLS84, Sho84, LV89a]. cellular [Fre89]. Chain [Blu84]. chains [Mih89]. channel [KWW89]. Characterization [Mul84, Pit84]. Characterizations [CK89, FL89]. Chervonenkis [LMR88b]. Chicago [IEE82]. chordal [Kle88]. chromatic [LV89b]. Circuits [BCH84, All89, Cai89, CL89a, LMN89]. circulation [GPT88]. class [PS89]. Classes [BBL^{+}84, Sie89]. closure [PT88]. Coin [Blu84]. Coins [BD84]. Combinatorial [CEG^{+}88, GPT88, Mih89]. Common [KMR84, KT88, MST88]. Communication [MR84, PU84, PS84, Pip84, Tiw84, AMS89, DF89, Kus89, RW89]. Communication-Time [PU84]. communications [LS88]. Comparative [Agg84]. comparison [AA88]. completeness [BSS88]. Complexity [BS84a, BBL^{+}84, Cha84, GS84, MSM84, PS84, Tiw84, AFEM89, Ajt88, CRDX88, CEG^{+}88, CL89b, DF89, EJ88, IL89a, KSS89, Ko89, Kus89, LV89a, LL89, LS88, MT89, MST89, MNR89, Pla89, RW89, Saf88]. Complexity-Preserving [Cha84]. Components [TV84]. Compression [GO84, HS84]. computation [ALGP89, BG89, BSS88, GSPT89, Lub88, Yao89]. Computational [Ko89, Tod89]. computations [Kos89b, MST88]. Computer [Ano84, IEE80, IEE81, IEE82, IEE83, IEE84, IEE85, IEE86, IEE87, IEE88, IEE89, Tiw84]. Computers [FS84]. Computing [BR89a, Cha84, KT88, KMR84, TV84, DS89a, FL89]. concepts [LV89a]. Conductance [Mih89]. Congruential [FKL84, Ste87]. conjecture [Sun89]. connectivity [KS89, KSS89, KP88]. consensus [BGP89]. Constant [LMN89, Cai89, Koc88]. constrained [HS89]. Construct [GGM84]. Constructing [Fri84]. Constructive [MRS88]. continued [Jus89]. Contraction [GGS84]. convergence [Hau89, Mih89]. convex [Cha89, Vai89a]. Coordinated [MW88]. Coordinating [KMS84]. Correlated [Blu84]. Cost [BS84b]. counterexamples [MT89]. cover [BRS89, BK88]. cryptographic [IN89]. Cryptographically [BM82, Ste87]. cryptography [IL89a]. Cryptosystems [GS84]. curves [CEG^{+}88]. cut [GH88, LR88]. cycle [DHK88].

Data [GO84, Mun84, BV89]. data-structure [BV89]. dataflow [Rus89]. Datalog [AG89]. Davenport [HS84]. Decidability [AH89a]. deciding [Ren88]. Decision [IT89, MSM84]. decomposable [DS89a]. decomposition [ALGP89]. degree [Her89]. Degrees [Hom84]. denominators [KT88]. dense [DHK88]. Depth [BCH84, Cai89, LMN89]. deque [Sun89]. Designing [IPK84]. deterministic [CF88, CW89, MNN89]. Diagnosability [Sul84]. Diameter [KMS84]. Dictionary [Mun84, MNR89]. dimension [Cla88, Hau89, LMR88b]. dimension-based [Hau89]. Dimensional [Vai84, Cha89]. dimensions [For89, Jus89]. directed [AF88]. disjoint [KS89, KMV89]. Dispersers [CW89]. distinctness [Yao88]. Distributed [IR84, Tiw84, UW84, ALGP89, BGP89]. distribution [LV89a]. distributions [LV89a]. Division [BCH84]. Divisor [KMR84, MST88]. divisors [KT88]. Do [FKL84]. Double [Mil89]. Down [Col84]. drawn [HLW88]. Dynamic [AS88, DKM^{+}88, IA84, AGR89, Awe88, EGG88, PT88]. Dynamically [DS89a].

Edges [Mis84]. Effect [KP88]. effects [Awe88]. Efficient [BRS89, GGS84, GX89, Her89, IN89, KS89, Kle88, Kos89a, Mis84, VV84, DK89, Jac89, KMV89]. Eigenvalues [AM84]. element [Yao88]. Elementary [Fri84]. elimination [DK89]. Embedding [Hea84]. embeddings [Fre89, Han89, MRS88]. Enclosure [CY84]. end [AMS89]. end-to-end [AMS89]. Ensemble [FG89]. environments [CM89]. equality [PS89]. equivalence [Tse89]. essential [IL89a]. Euclidean [SV84b]. Evaluating [BS84b]. evaluation [GK89]. evaluations [KT88]. Every [CS84]. existence [GKL88]. existential [PS89, Ren88]. Expanders [AM84, LM89, Mih89]. expression [GK89]. expressiveness [AH89a].

factor [Koc88]. factoring [Ron89].

factorization [KT88]. factors [DLMV88]. fails [PS89]. fairness [MPS88]. Fast [BLS88, Ebe84, MS89, Vai84, AS88, DH89, IL89b, LM89, Mul88, Sie89, Vai89b]. Faster [Col84, Ren88]. Fault [Sul84]. faults [LM89]. faulty [BG89]. feasible [CK89]. feedback [Awe88]. Fibonacci [FT84]. fields [Ron89, Sho88]. Find [Mis84]. Finding [TV84, KS89, KMV89, Sho88]. Finite [BS84b, Blu84, GW84, AF88, BR89a, CK89, CY88, DS89b, FL89, Jia89, Ron89, Sho88]. finite-basis [FL89]. finite-state [CY88]. first [AH89a, AG89]. first-order [AH89a, AG89]. Fishspear [FP84]. fixed [AFEM89, Mul89]. Flipping [BD84]. Flips [Blu84]. Florida [IEE84]. Flow [MN89, CH89, LR88]. Formulae [Fri84]. Foundations [Ano84, IEE80, IEE81, IEE82, IEE83, IEE84, IEE85, IEE86, IEE87, IEE88, IEE89]. Fourier [LMN89]. fraction [Jus89]. Free [Cha84]. Full [Mul84, Rus89]. Fully [Ong88, PT88]. functionals [CK89]. Functions [BS84b, GGM84, TV84, BSS88, DMW88, DS89a, HLW88, IL89a, KKL88, Ko89, LS88, Sie89, Vai89a].

Galois [Ron89]. game [MNR89]. general [Mil89]. Generalized [HS84, GPT88]. Generalizing [Hau89, Jus89]. Generate [BM82]. Generating [Bro89, SV84a]. Generation [VV84]. Generators [FKL84, GKL88, KGY89, Ste87]. Genus [Mal88]. geometric [Han89]. geometry [BRS89, CF88, Mil89]. given [KT88]. Gödel [PS89]. Good [BCLS84]. Graph [BCLS84, GGS84, LV89b, Mis84, CFI89, DK89, Fre89, MRS88, RR89]. graphic [GX89]. Graphs [Hea84, KMS84, SV84b, AF88, DLMV88, DHK88, Jac89, KS89, Kle88, Mal88, MN89]. Greatest [KMR84, KT88, MST88]. Group [BS84a]. Groups [KMS84, BLS88, BR89a, FMS^{+}89, Ron89]. grow [BC88].

Hamiltonian [DHK88]. harder [AF88]. Hardness [NW88]. Hardware [MR84]. hash [Sie89]. hashing [DKM^{+}88]. Heaps [FT84]. help [Cai89]. hidden [OS89]. high [Sie89]. homeomorph [KMV89]. Homogeneous [Lev88]. Honest [Hom84].

identification [CFI89]. Illinois [IEE82]. Implementation [GGS84]. Implicit [Mun84]. Improved [FT84]. Inclusion [CY84]. increase [Koc88]. Increasing [Koc88]. Incremental [BT89]. independence [BR89b]. Independent [Blu84, GX89]. Inference [Pit84]. Infinite [BS84b]. influence [KKL88]. inputs [Yao89]. insensitive [IL89b]. integer [MST88, Yao89]. Integers [Kam84, KMR84]. Interactive [GO84, CL89b]. Interior [GSPT89]. Interior-point [GSPT89]. Intersecting [DE84, CE88, Cha89]. Intersection [IA84]. intersections [Mil89]. invariants [Lev88]. inverse [Dic89]. irreducible [BR89a, Sho88]. Island [IEE84].

Key [GS84]. kinodynamic [CRDX88].

Klee [OY88]. Knowledge [FHV84, Kil88, KMO89]. Kuratowski [KMV89].

Lambda [Mit84, Mul84, Ong88]. Large [MR84, DLMV88]. Lattices [LS88]. law [PS89]. layer [KWW89]. Layout [FS84]. lazy [Ong88]. learnability [LMR88b, LMN89, Sch89]. Learning [DMW88, Gas88, GRS89, BRS89, FMS^{+}89, LV89a, MT89, Tse89]. Limited [Pip84]. Limits [MR84]. Line [Sho84, CE88, Mil89]. Linear [FKL84, Kom84, SG84, Cla88, GX89, Ste87, Vai89b]. linkless [MRS88]. locality [ALGP89]. locally [Kre89]. location [PT88]. Log [BCH84, Kil88]. log-space [Kil88]. Logarithmic [TV84]. logic [AG89, AH89b]. logics [AH89a, EJ88]. Loose [CS84]. Lower [Cai89, GW84, KGY89, MST88, Ng89, Tiw84, Yao89, AGR89, Bsh88, CFI89, DKM^{+}88].

machine [GK89]. Machines [Agg84, GW84, IPK84, MR84, BSS88, LL89]. maintenance [For89]. majority [BG89, LW89]. management [BLS88]. Many [BD84]. Mappings [Cha84]. Markov [Blu84, Mih89]. marriage [Ng89]. matching [GPV88, Kos89a, MS89]. Matrix [BS84a, Ebe84, Bsh88, Vai89b]. Matroids [SG84, GX89]. max [LR88]. max-flow [LR88]. maxima [DS89a]. maximum [CH89]. maximum-flow [CH89]. measure [OY88]. Measures [GS84, Lev88]. memories [Her89]. Memory [UW84, KP88]. MEO [DK89]. method [MNN89]. methods [GSPT89]. metric [Hau89]. might [LM89]. min [LR88]. min-cut [LR88]. Minimal [Hom84, DK89]. minimizing [Vai89a]. Minimum [KMO89, Vai84]. minors [MRS88]. Möbius [LS88]. Model [FHV84, Hau89]. Model-Theoretic [FHV84]. Models [Mit84, IL89b, KP88, Ong88]. Monotone [Fri84, AP88]. Motion [KMS84, FG89]. Multi [IR84]. Multi-Tree [IR84]. multibutterflies [LM89]. multicommodity [LR88]. multidimensional [AP88]. Multiparty [BG89, DF89]. multiple [MN89]. Multiplication [Kam84, Bsh88, Vai89b]. Myhill [FL89]. Myhill-Nerode [FL89].

Nashville [IEE81]. NC [BRS89, BR89b]. Near [Yao88]. Near-optimal [Yao88]. Nerode [FL89]. nested [Lan89]. Network [ALGP89, FT84, Awe88, Koc88, Pla89]. Networks [Col84, IR84, Tiw84, AGR89, AS88, Her89, Jia89, Rus89]. nondeterministic [Rus89]. Nonexpressibility [MPS88]. Nonlinearity [HS84]. nonuniform [Jia89]. North [IEE89]. note [All89]. Notes [AP88]. Novel [FS84]. November [IEE82, IEE83, IEE89]. NP [BSS88]. Number [VV84, CFI89, KGY89, PSS89]. numbers [BSS88, LV89b]. numerators [KT88].

Objects [DE84]. oblivious [CK88]. obstructions [Mul89]. Obtain [Col84]. October [IEE80, IEE81, IEE84, IEE85, IEE86, IEE87, IEE88, IEE89]. On-Line [Sho84]. One [IL89a]. One-way [IL89a]. Optimal [DHK88, BGP89, CFI89, CE88, Cha89, Kre89, KWW89, RR89, Yao88]. Optimization [FT84]. Oracles [BBL^{+}84]. Order [Mit84, AH89a, AG89, PS89]. ordering [DK89]. orders [GRS89]. Oregon [IEE85]. Other [CY84]. Output [OS89]. Output-sensitive [OS89].

P [Tod89]. PAC [Hau89]. packet [LMR88a]. Packing [Sho84]. pagenumber [Mal88]. Pages [Hea84]. pairs [Fre89]. papers [IEE81]. Paradoxical [GMR84]. Parallel [AA88, Ebe84, KMR84, MR84, Pip84, TV84, vzG84, BV89, DHK88, DK89, GPV88, GSPT89, GK89, Han89, KS89, KMV89, KSS89, Kle88, Lub88, MNN89, RR89]. parameter [AFEM89]. Parity [SG84]. Park [IEE89]. partition [Mul88]. Path [HS84, SG84]. Paths [SV84b, Fre89, KS89, KMV89]. pattern [Kos89a]. Pebble [KMS84]. Peeling [CY84]. penalty [Lub88]. Rabin [ACGS84]. perfect [DKM^{+}88]. performance [Koc88, Sie89]. permanents [DLMV88]. Permutation [KMS84, BLS88, FMS^{+}89, vzG89]. pigeonhole [Ajt88]. Pipelining [Kos89b]. Plains [IEE88]. Planar [Hea84, MN89, Mul88, PSS89, PT88]. planarity [BT89, RR89]. plane [CE88, Han89]. Planning [FMS^{+}89, CRDX88, FG89]. Pockets [BD84]. point [For89, GSPT89, PT88]. pointer [GK89]. pointers [BAG88]. points [HLW88, MS89]. Polygon [CY84, MS89]. polyhedra [Cha89]. Polylog [Mun84]. Polymorphic [FS84]. Polynomial [AMS89, CY84, Ebe84, Fri84, GH88, Hom84, Sul84, Dic89, IT89, Lev88, MW88]. Polynomials [Kam84, KT88, Ron89, Sho88, vzG89]. Polytopes [DLMV88]. Portland [IEE85]. Potato [CY84]. Potato-Peeling [CY84]. Power [IL89b, All89, DS89b, Tod89]. Powering [vzG84]. PP [Tod89]. practical [LM89]. Precision [BS84b, Mil89]. Predicting [HLW88]. prediction [DMW88]. Preliminary [FHV84]. presence [Cai89]. Preserving [Cha84]. principle [Ajt88]. Priority [FP84]. Privacy [Kus89]. Probabilistic [GW84, PS84, Pit84, RW89, CY88, DMW88, DS89b, MNN89, Tse89]. probability [AH89a]. Problem [GMR84, Mun84, SG84, DHK88, GPT88, GH88, KMV89, KSS89, MNR89, Ng89, OY88]. Problems [BS84a, BCH84, CY84, AFEM89, AA88, Fre89, GPV88, HS89, Han89, IT89, LR88]. processing [Han89]. Processor [KMV89, Lub88]. processors [Kos89b].

Produce [FKL84]. products [LV89b]. programming [Cla88, EGG88, Vai89b]. programs [CL89a, CY88, EJ88]. proofs [CL89b, KMO89]. properties [CY88]. protocols [Awe88]. provably [IN89]. Pseudo [BM82, VV84]. Pseudo-Random [BM82, VV84]. pseudorandom [GKL88, KGY89]. PSPACE [Ren88]. Public [GS84]. Public-Key [GS84]. Pyramid [Agg84].

Quasi [SV84a]. Quasi-Random [SV84a]. queries [Gas88]. Queue [FP84].

radicals [Lan89]. Ramsey [MSM84]. Randolli [GGM84]. Random [BM82, BD84, FKL84, SV84a, VV84, Bro89, CF88, CW89, IZ89]. Randomized [AS89, CH89]. randomly [HLW88]. randomness [Lub88, NW88]. Rational [BS84b]. Reachability [AF88]. real [BSS88, Ko89]. really [AH89b]. reals [Ren88]. Recursive [BV89, BSS88]. recycle [IZ89]. Reducibilities [Hom84]. Related [Agg84, BCH84, DH89, GPV88, MNR89]. relation [Mul89]. relations [GRS89, RW89]. Reliability [IR84]. removal [OS89]. Removing [Lub88]. Report [FHV84]. representations [BR89a]. Research [IEE89]. resource [KMO89]. Results [LMR88b, Hau89, MRS88]. reversal [LL89]. Ring [Kam84]. River [CS84]. root [MST89]. roots [Ko89]. round [MW88]. rounded [Mil89]. Routing [CS84, AGR89, KWW89, LMR88a, LM89]. RSA [ACGS84]. RSA/Rabin [ACGS84]. Runs [Mun84].

sample [Hau89]. sampling [CF88]. scanning [Sun89]. scheduling [HS89]. Scheme [HS84]. schemes [AGR89, HS89, IN89]. Schinzel [HS84]. Science [IEE80, IEE81, IEE82, IEE83, IEE84, IEE85, IEE86, IEE87, IEE88, IEE89, Ano84]. Search [IA84, AS89, IT89]. Searching [DE84, AP88]. Second [Mit84, PS89]. Second-Order [Mit84]. Secret [Ste87]. Secure [ACGS84, VV84, IN89, Ste87]. security [CK88]. Segment [IA84, Mil89]. segments [CE88]. selection [Pla89]. Semantic [Mit84, Mul84]. sensitive [OS89]. sentences [PS89]. separation [KT88]. Sequences [BM82, FKL84, HS84, SV84a]. Sequential [IPK84]. set [BRS89, For89, GK89]. sets [PSS89, Vai89a]. Seven [Hea84]. Share [UW84]. shared [Her89]. Shortest [SV84b, Fre89]. signaling [MPS88]. Signature [GMR84]. simple [LV89a]. Simplification [Lan89]. Simulating [BR89b]. simulations [CL89a, Her89]. Singer [IEE84]. sinks [MN89]. Size [Fri84, Hau89, Koc88]. Slightly [SV84a]. Slightly-Random [SV84a]. Slowing [Col84]. small [Cla88, Her89]. Solution [CY84, GMR84]. solutions [Kre89]. Solvability [CM89]. solving [Fre89]. Some [Sho84]. Sorting [Col84, GK89]. Source [Blu84]. Sources [SV84a, CW89, MN89]. Space [DE84, Jac89, Vai84, CL89b, Kil88, Sie89, Yao88]. Space-efficient [Jac89]. Spanning [Kom84, Vai84, Bro89]. Sparse [BBL^{+}84]. Speed [MR84]. Speeding [EGG88, Vai89b]. Speeding-up [Vai89b]. spherical [DH89]. square [MST89]. Stable [For89, Ng89]. stacks [FG88]. State [Blu84, GW84, CY88, DS89b]. static [AS88, Jac89]. strength [Sch89]. Strong [BM82]. Structure [Kre89, Mun84, BV89]. structures [PT88]. Study [Agg84]. subgraph [KSS89]. Sublinear [GPV88, KMR84]. Sublinear-time [GPV88]. Subquadratic [CL89a]. subset [IN89]. sum [IN89]. super [IT89]. super-polynomial [IT89]. Superconcentrators [AM84]. surface [OS89]. surfaces [CEG^{+}88]. Symmetric [Fri84]. Symposium [Ano84, IEE80, IEE81, IEE82, IEE83, IEE84, IEE85, IEE86, IEE87, IEE88, IEE89]. synchronization [Jia89]. Syracuse [IEE80]. System [UW84]. Systolic [FS84, IPK84].

Take [BC88]. technique [Mil89]. techniques [PT88]. temporal [AH89b, CY88]. Tennessee [IEE81]. Testing [vzG89, BT89, KS89]. th [Fri84]. Their [FT84, KT88, Sie89]. Theorem [MSM84, FL89, LR88, Sun89]. Theoretic [FHV84]. theory [BSS88, LV89a, Ren88]. thinness [IL89b]. Three [FG88, Cha89, KWW89]. three-dimensional [Cha89]. three-layer [KWW89]. threshold [All89]. Time [Mun84, PU84, Sul84, TV84, BK88, Dic89, GPV88, IT89, Lev88, MW88, Sie89, Yao88]. time-space [Sie89, Yao88]. Toronto [IEE86]. total [GRS89]. Tractable [BS84b]. Tradeoff [PU84, Sie89, Yao88]. transfer [CK88]. transform [LMN89]. transforms [DH89]. transitive [PT88]. traversal [MW88]. treatment [Mih89]. Tree [Agg84, Cha84, IR84, TV84, BV89, BC88, EJ88, Kos89a, Kos89b]. Trees [Kom84, MSM84, Vai84, AS89, Bro89, FG89, Jac89, Yao89]. Triangle [IEE89]. triangulations [For89]. Tucson [IEE83]. Turing [LL89]. turns [Sun89]. Twists [Sun89]. Two [KMR84, For89, KMV89]. type [CK89]. Typed [Mul84].

Unbiased [Blu84]. Unbounded [MR84]. Undirected [Mis84, AF88]. Uniform [BBL^{+}84, Hau89, LR88]. Uniformly [BD84]. Universal [LMR88a, BSS88, LV89a, Sie89]. Upper [AGR89, DKM^{+}88, OY88, PSS89]. use [CF88, FL89]. Uses [FT84]. Using [Fre89, GGS84, IPK84, CK88, Mil89, Vai89b].

Values [BD84]. Vapnik [LMR88b].

Variables [Fri84, CFI89, KKL88]. variants [Ng89]. Vegas [Cla88]. Verification [Kom84]. verifiers [Kil88]. Verifying [CY88]. versus [BAG88, IT89]. Very [Ebe84]. Via [Cha84, Gas88]. view [CF88]. viewpoint [Mul89]. VLSI [FS84, IL89b]. vs [AG89, NW88].

walk [BC88]. Way [CS84, DS89b, IL89a]. weak [CW89, Sch89]. weakened [CK88]. weighted [LW89]. Which [CS84]. White [IEE88]. wire [IL89b]. wise [BR89b]. without [Lub88].

X [Agg84]. X-Tree [Agg84].

yields [MNN89]. York [IEE80, IEE88].

Zero [Kil88, KMO89]. Zero-knowledge [Kil88].

References

Alon:1988:PCA

[AA88] N. Alon and Y. Azar. Parallel comparison algorithms for approximation problems. In IEEE [IEE88], pages 194–203. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Alexi:1984:RRB

[ACGS84] W. Alexi, B. Chor, O. Goldreich, and C. P. Schnorr. RSA/ Rabin bits are 1/2 + 1Poly(log N) secure. In IEEE [IEE84], pages 449–457. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Ajtai:1988:RHD

[AF88] M. Ajtai and R. Fagin. Reachability is harder for directed than for undirected finite graphs. In IEEE [IEE88], pages 358–367. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Abrahamson:1989:CFP

[AFEM89] K. R. Abrahamson, M. R. Fellows, J. A. Ellis, and M. E. Mata. On the complexity of fixed parameter problems. In IEEE [IEE89], pages 210–215. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Ajtai:1989:DVF

[AG89] M. Ajtai and Y. Gurevich. Datalog vs. first-order logic. In IEEE [IEE89], pages 142–147. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Aggarwal:1984:CSX

[Agg84] A. Aggarwal. A comparative study of X-tree, pyramid and related machines. In IEEE [IEE84], pages 89–99. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Afek:1989:ULB

[AGR89] Y. Afek, E. Gafni, and M. Ricklin. Upper and lower bounds for routing schemes in dynamic networks. In IEEE [IEE89], pages 370–375. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Abadi:1989:DEF

[AH89a] M. Abadi and J. Y. Halpern. Decidability and expressiveness for first-order logics of probability. In IEEE [IEE89], pages 148–153. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Alur:1989:RTL

[AH89b] R. Alur and T. A. Henzinger. A really temporal logic. In IEEE [IEE89], pages 164–169. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Ajtai:1988:CPP

[Ajt88] M. Ajtai. The complexity of the pigeonhole principle. In IEEE [IEE88], pages 346–355. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Awerbuch:1989:NDL

[ALGP89] B. Awerbuch, M. Luby, A. V.

Goldberg, and S. A. Plotkin. Network decomposition and locality in distributed computation. In IEEE [IEE89], pages 364–369. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Allender:1989:NPT

[All89] E. Allender. A note on the power of threshold circuits. In IEEE [IEE89], pages 580–584. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Alon:1984:EES

[AM84] N. Alon and V. D. Milman. Eigenvalues, expanders and superconcentrators. In IEEE [IEE84], pages 320–322. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Awerbuch:1989:PEE

[AMS89] B. Awerbuch, Y. Mansour, and N. Shavit. Polynomial end-to-end communication. In IEEE [IEE89], pages 358–363. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Anonymous:1984:ASF

[Ano84] Anonymous. 25th Annual Symposium on Foundations of Computer Science. In IEEE [IEE84], pages i–xii. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Aggarwal:1988:NSM

[AP88] A. Aggarwal and J. Park. Notes on searching in multidimensional monotone arrays. In IEEE [IEE88], pages 497–512. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Awerbuch:1988:DNF

[AS88] B. Awerbuch and M. Sipser. Dynamic networks are as fast as static networks. In IEEE [IEE88], pages 206–219. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Aragon:1989:RST

[AS89] C. R. Aragon and R. G. Seidel.

Randomized search trees. In IEEE [IEE89], pages 540–545. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Awerbuch:1988:EFD

[Awe88] B. Awerbuch. On the effects of feedback in dynamic network protocols. In IEEE [IEE88], pages 231–245. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Ben-Amram:1988:PVA

[BAG88] A. M. Ben-Amram and Z. Galil. On pointers versus addresses. In IEEE [IEE88], pages 532–538. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Balcazar:1984:SOU

[BBL^{+}84] J. Balcazar, R. Book, T. Long, U. Schoning, and A. Selman. Sparse oracles and uniform complexity classes. In IEEE [IEE84], pages 308–311. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Bhatt:1988:TWG

[BC88] S. Bhatt and Jin-Yi Cai. Take a walk, grow a tree. In IEEE [IEE88], pages 469–478. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Beame:1984:LDC

[BCH84] P. W. Beame, S. A. Cook, and H. J. Hoover. Log depth circuits for division and related problems. In IEEE [IEE84], pages 1–6. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Bui:1984:GBA

[BCLS84] Thang Bui, S. Chaudhuri, T. Leighton, and M. Sipser. Graph bisection algorithms with good average case behavior. In IEEE [IEE84], pages 181–192. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Broder:1984:FCM

[BD84] A. Z. Broder and D. Dolev.

Flipping coins in many pockets (Byzantine agreement on uniformly random values). In IEEE [IEE84], pages 157–170. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Beaver:1989:MCF

[BG89] D. Beaver and S. Goldwasser. Multiparty computation with faulty majority. In IEEE [IEE89], pages 468–473. CODEN ASF-PDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Berman:1989:TOD

[BGP89] P. Berman, J. A. Garay, and K. J. Perry. Towards optimal distributed consensus. In IEEE [IEE89], pages 410–415. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Broder:1988:BCT

[BK88] A. Z. Broder and A. R. Karlin. Bounds on the cover time. In IEEE [IEE88], pages 479–487. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Babai:1988:FMP

[BLS88] L. Babai, E. M. Luks, and A. Seress. Fast management of permutation groups. In IEEE [IEE88], pages 272–282. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Blum:1984:IUC

[Blu84] M. Blum. Independent unbiased coin flips from a correlated biased source: A finite state Markov chain. In IEEE [IEE84], pages 425–433. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Blum:1982:HGC

[BM82] Manuel Blum and Silvio Micali.

How to generate cryptographically strong sequences of pseudorandom bits. In IEEE [IEE82], pages 112–117. CODEN ASF-PDV. ISBN ???? ISSN 0272-5428. LCCN QA76.6 .S95 1982. IEEE catalog no. 82CH1806-9. IEEE Computer Society order no. 440.

Babai:1989:CIR

[BR89a] L. Babai and L. Ronyai. Computing irreducible representations of finite groups. In IEEE [IEE89], pages 93–98. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Berger:1989:SWI

[BR89b] B. Berger and J. Rompel. Simulating (log^{c} n)-wise independence in NC. In IEEE [IEE89], pages 2–7. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Broder:1989:GRS

[Bro89]  A. Broder.  Generating ran-

dom spanning trees. In IEEE [IEE89], pages 442–447. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Berger:1989:ENA

[BRS89] B. Berger, J. Rompel, and P. W. Shor. Efficient NC algorithms for set cover with applications to learning and geometry. In IEEE [IEE89], pages 54–59. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Babai:1984:CMG

[BS84a] L. Babai and E. Szemeredi. On the complexity of matrix group problems I. In IEEE [IEE84], pages 229–240. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Blum:1984:ERF

[BS84b] L. Blum and M. Shub. Evaluating rational functions: Infinite precision is finite cost and tractable on average. In IEEE [IEE84], pages 261–267. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Bshouty:1988:LBM

[Bsh88] N. H. Bshouty. A lower bound for matrix multiplication. In IEEE [IEE88], pages 64–67. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Blum:1988:TCR

[BSS88] L. Blum, M. Shub, and S. Smale. On a theory of computation over the real numbers; NP completeness, recursive functions and universal machines. In IEEE [IEE88], pages 387–397. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Battista:1989:IPT

[BT89] G. Di Battista and R. Tamassia. Incremental planarity testing. In IEEE [IEE89], pages 436– 441. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Berkman:1989:RTP

[BV89] O. Berkman and U. Vishkin. Recursive *-tree parallel datastructure. In IEEE [IEE89], pages 196–202. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Cai:1989:LBC

[Cai89] J.-Y. Cai. Lower bounds for constant depth circuits in the presence of help bits. In IEEE [IEE89], pages 532–537. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Chazelle:1988:OAI

[CE88] B. Chazelle and H. Edelsbrunner.

An optimal algorithm for intersecting line segments in the plane. In IEEE [IEE88], pages 590–600. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Clarkson:1988:CCB

[CEG^{+}88] K. L. Clarkson, H. Edelsbrunner, L. J. Guibas, M. Sharir, and E. Welzl. Combinatorial complexity bounds for arrangements of curves and surfaces. In IEEE [IEE88], pages 568–579. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Chazelle:1988:DVR

[CF88] B. Chazelle and J. Friedman. A deterministic view of random sampling and its use in geometry. In IEEE [IEE88], pages 539–549. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Cai:1989:OLB

[CFI89] J.-Y. Cai, M. Furer, and N. Immerman. An optimal lower bound on the number of variables for graph identification. In IEEE [IEE89], pages 612–617. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Cheriyan:1989:RMF

[CH89] J. Cheriyan and T. Hagerup. A randomized maximum-flow algorithm. In IEEE [IEE89], pages 118–123. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Chazelle:1984:CFT

[Cha84] B. Chazelle. Computing on a free tree via complexity-preserving mappings. In IEEE [IEE84], pages 358–368. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Chazelle:1989:OAI

[Cha89] B. Chazelle. An optimal algorithm for intersecting three-dimensional convex polyhedra. In IEEE [IEE89], pages 586–591. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Crepeau:1988:AOT

[CK88] C. Crepeau and J. Kilian. Achieving oblivious transfer using weakened security assumptions. In IEEE [IEE88], pages 42–52. CO-DEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Cook:1989:CBF

[CK89] S. A. Cook and B. M. Kapron.

Characterizations of the basic feasible functionals of finite type.

In IEEE [IEE89], pages 154–159. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Cai:1989:SSC

[CL89a] J.-Y. Cai and R. J. Lipton. Subquadratic simulations of circuits by branching programs. In IEEE [IEE89], pages 568–573. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Condon:1989:CSB

[CL89b] A. Condon and R. J. Lipton. On the complexity of space bounded interactive proofs. In IEEE [IEE89], pages 462–467. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Clarkson:1988:VAL

[Cla88] K. L. Clarkson. A Las Vegas algorithm for linear programming when the dimension is small. In IEEE [IEE88], pages 452–456. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Chor:1989:SAE

[CM89] B. Chor and L. Moscovici. Solvability in asynchronous environments. In IEEE [IEE89], pages 422–427. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Cole:1984:SSN

[Col84] R. Cole. Slowing down sorting networks to obtain faster sorting algorithm. In IEEE [IEE84], pages 255–260. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Canny:1988:CKP

[CRDX88] J. Canny, J. Reif, B. Donald, and P. Xavier. On the complexity of kinodynamic planning. In IEEE [IEE88], pages 306–316. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Cole:1984:RRE

[CS84] R. Cole and A. Siegel. River routing every which way, but loose. In IEEE [IEE84], pages 65–73. CO-DEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Cohen:1989:DDA

[CW89] A. Cohen and A. Wigderson. Dispersers, deterministic amplification, and weak random sources. In IEEE [IEE89], pages 14–19. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Chang:1984:PSP

[CY84] J. S. Chang and C. K. Yap. A polynomial solution for potatopeeling and other polygon inclusion and enclosure problems. In IEEE [IEE84], pages 408–416. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Courcoubetis:1988:VTP

[CY88] C. Courcoubetis and M. Yannakakis. Verifying temporal properties of finite-state probabilistic programs. In IEEE [IEE88], pages 338–345. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Dobkin:1984:SSI

[DE84] D. P. Dobkin and H. Edelsbrunner. Space searching for intersecting objects. In IEEE [IEE84], pages 387–392. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Dolev:1989:MCC

[DF89] D. Dolev and T. Feder. Multiparty communication complexity. In IEEE [IEE89], pages 428– 433. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Driscoll:1989:AFA

[DH89] J. R. Driscoll and D. M. Healy, Jr.

Asymptotically fast algorithms for spherical and related transforms.

In IEEE [IEE89], pages 344–349. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Dahlhaus:1988:OPA

[DHK88] E. Dahlhaus, P. Hajnal, and M. Karpinski. Optimal parallel algorithm for the Hamiltonian cycle problem on dense graphs. In IEEE [IEE88], pages 186–193. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Dickerson:1989:IAP

[Dic89] M. Dickerson. The inverse of an automorphism in polynomial time. In IEEE [IEE89], pages 82– 87. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Dahlhaus:1989:EPA

[DK89] E. Dahlhaus and M. Karpinski. An efficient parallel algorithm for the minimal elimination ordering (MEO) of an arbitrary graph. In IEEE [IEE89], pages 454–459. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Dietzfelbinger:1988:DPH

[DKM^{+}88] M. Dietzfelbinger, A. Karlin, K. Mehlhorn, F. M. auf der Heide, H. Rohnert, and R. E. Tarjan. Dynamic perfect hashing: upper and lower bounds. In IEEE [IEE88], pages 524–531. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Dagum:1988:PPG

[DLMV88] P. Dagum, M. Luby, M. Mihail, and U. Vazirani. Polytopes, permanents and graphs with large factors. In IEEE [IEE88], pages 412–421. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

DeSantis:1988:LPP

[DMW88] A. DeSantis, G. Markowsky, and M. N. Wegman. Learning probabilistic prediction functions. In IEEE [IEE88], pages 110–119. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Dobkin:1989:DCM

[DS89a] D. Dobkin and S. Suri. Dynamically computing the maxima of decomposable functions, with applications. In IEEE [IEE89], pages 488–493. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Dwork:1989:PWP

[DS89b] C. Dwork and L. Stockmeyer. On the power of 2-way probabilistic finite state automata. In IEEE [IEE89], pages 480–485. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Eberly:1984:VFP

[Ebe84] W. Eberly. Very fast parallel matrix and polynomial arithmetic. In IEEE [IEE84], pages 21–30. CO-DEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Eppstein:1988:SDP

[EGG88] D. Eppstein, Z. Galil, and R. Giancarlo. Speeding up dynamic programming. In IEEE [IEE88], pages 488–496. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Emerson:1988:CTA

[EJ88] E. A. Emerson and C. S. Jutla.

The complexity of tree automata and logics of programs. In IEEE [IEE88], pages 328–337. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Fredman:1988:TS

[FG88] M. L. Fredman and D. L. Goldsmith. Three stacks. In IEEE [IEE88], pages 514–523. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Frederickson:1989:EMP

[FG89] G. N. Frederickson and D. J. Guan. Ensemble motion planning in trees. In IEEE [IEE89], pages 66–71. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Fagin:1984:MTA

[FHV84] R. Fagin, J. Y. Halpern, and M. Y. Vardi. A model-theoretic analysis of knowledge: Preliminary report. In IEEE [IEE84], pages 268–278. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Frieze:1984:LCG

[FKL84] A. M. Frieze, R. Kannan, and J. C. Lagarias. Linear congruential generators do not produce random sequences. In IEEE [IEE84], pages 480–484. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Fellows:1989:AMN

[FL89] M. R. Fellows and M. A. Langston. An analogue of the Myhill-Nerode theorem and its use in computing finite-basis characterizations. In IEEE [IEE89], pages 520–525. CODEN ASF-PDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Fiat:1989:PLP

[FMS^{+}89] A. Fiat, S. Moses, A. Shamir, I. Shimshoni, and G. Tardos. Planning and learning in permutation groups. In IEEE [IEE89], pages 274–279. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Fortune:1989:SMP

[For89] S. Fortune. Stable maintenance of point set triangulations in two dimensions. In IEEE [IEE89], pages 494–499. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Fischer:1984:FPQ

[FP84] M. J. Fischer and M. S. Paterson. Fishspear: A priority queue algorithm. In IEEE [IEE84], pages 375–386. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Frederickson:1989:UCG

[Fre89] G. N. Frederickson. Using cellular graph embeddings in solving all pairs shortest paths problems. In IEEE [IEE89], pages 448– 453. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Friedman:1984:CSM

[Fri84]  J. Friedman.  Constructing

O(n log n) size monotone formulae for the k-th elementary symmetric polynomial of n Boolean variables. In IEEE [IEE84], pages 506–515. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Fiat:1984:PAN

[FS84] A. Fiat and A. Shamir. Polymorphic arrays: A novel VLSI layout for systolic computers. In IEEE [IEE84], pages 37–45. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Fredman:1984:FHT

[FT84] M. L. Fredman and R. E. Tarjan. Fibonacci heaps and their uses in improved network optimization algorithms. In IEEE [IEE84], pages 338–346. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Gasarch:1988:LQ

[Gas88] W. I. Gasarch. Learning via queries. In IEEE [IEE88], pages 130–137. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Goldreich:1984:HCR

[GGM84] O. Goldreich, S. Goldwasser, and S. Micali. How to construct Randolli functions. In IEEE [IEE84], pages 464–479. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Gabow:1984:EIG

[GGS84] H. N. Gabow, Z. Galil, and T. H.

Spencer. Efficient implementation of graph algorithms using contraction. In IEEE [IEE84], pages 347–357. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Goldschmidt:1988:PAK

[GH88] O. Goldschmidt and D. S. Hochbaum. Polynomial algorithm for the k-cut problem. In IEEE [IEE88], pages 444–451. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Goodrich:1989:SPP

[GK89] M. T. Goodrich and S. R. Kosaraju. Sorting on a parallel pointer machine with applications to set expression evaluation. In IEEE [IEE89], pages 190– 195. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Goldreich:1988:EPG

[GKL88] O. Goldreich, H. Krawczyk, and M. Luby. On the existence of pseudorandom generators. In IEEE [IEE88], pages 12–24. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Goldwasser:1984:SSS

[GMR84] S. Goldwasser, S. Micali, and R. L.

Rivest. A “paradoxical” solution to the signature problem. In IEEE [IEE84], pages 441–448. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Gamal:1984:IDC

[GO84] A. El Gamal and A. Orlitsky. Interactive data compression. In IEEE [IEE84], pages 100–108. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Goldberg:1988:CAG

[GPT88] A. V. Goldberg, S. A. Plotkin, and E. Tardos. Combinatorial algorithms for the generalized circulation problem. In IEEE [IEE88], pages 432–443. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Goldberg:1988:STP

[GPV88] A. V. Goldberg, S. A. Plotkin, and P. M. Vaidya. Sublinear-time parallel algorithms for matching and related problems. In IEEE [IEE88], pages 174–185. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Goldman:1989:LBR

[GRS89] S. A. Goldman, R. L. Rivest, and R. E. Schapire. Learning binary relations and total orders. In IEEE [IEE89], pages 46–51. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Grollmann:1984:CMP

[GS84] J. Grollmann and A. L. Selman. Complexity measures for publickey cryptosystems. In IEEE [IEE84], pages 495–503. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Goldberg:1989:IPM

[GSPT89] A. V. Goldberg, D. B. Shmoys, S. A. Plotkin, and E. Tardos.

Interior-point methods in parallel computation. In IEEE [IEE89], pages 350–355. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Greenberg:1984:LBP

[GW84] A. G. Greenberg and A. Weiss.

A lower bound for probabilistic algorithms for finite state machines. In IEEE [IEE84], pages 323–331. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Gabow:1989:EAI

[GX89] H. N. Gabow and Y. Xu. Efficient algorithms for independent assignment on graphic and linear matroids. In IEEE [IEE89], pages 106–111. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Hansen:1989:AAG

[Han89] M. D. Hansen. Approximation algorithms for geometric embeddings in the plane with applications to parallel processing problems. In IEEE [IEE89], pages 604– 609. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Haussler:1989:GPM

[Hau89] D. Haussler. Generalizing the PAC model: sample size bounds from metric dimension-based uniform convergence results. In IEEE [IEE89], pages 40–45. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Heath:1984:EPG

[Hea84] L. Heath. Embedding planar graphs in seven pages. In IEEE [IEE84], pages 74–83. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Herley:1989:ESS

[Her89] K. T. Herley. Efficient simulations of small shared memories on bounded degree networks.

In IEEE [IEE89], pages 390–395. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Haussler:1988:PFR

[HLW88] D. Haussler, N. Littlestone, and M. K. Warmuth. Predicting (0, 1)- functions on randomly drawn points. In IEEE [IEE88], pages 100–109. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Homer:1984:MDH

[Hom84] S. Homer. Minimal degrees for honest polynomial reducibilities. In IEEE [IEE84], pages 300–307. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Hart:1984:NDS

[HS84] S. Hart and M. Sharir. Nonlinearity of Davenport–Schinzel sequences and of a generalized path compression scheme. In IEEE [IEE84], pages 313–319. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Hall:1989:ASC

[HS89] L. A. Hall and D. B. Shmoys. Approximation schemes for constrained scheduling problems. In IEEE [IEE89], pages 134–139. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Imai:1984:DSI

[IA84] H. Imai and T. Asano. Dynamic segment intersection search with applications. In IEEE [IEE84], pages 393–402. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

IEEE:1980:ASF

[IEE80] IEEE, editor. 21st annual Symposium on Foundations of Computer Science: October 13–15, 1980, Syracuse, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1980. CODEN ASFPDV. ISBN ???? ISSN 0272-5428. LCCN QA76.6 .S95 1980; TK7885.A1 S92 1980. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 80CH1498-5.

IEEE:1981:ASF

[IEE81] IEEE, editor. 22nd Annual Symposium on Foundations of Computer Science: October 28–30, 1981, [Nashville, Tennessee: papers]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. CODEN ASF-PDV. ISBN ???? ISSN 0272-5428. LCCN TK7885.A1 S92 1981; QA76.6 .S95 1981. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 81CH1695-6.

IEEE:1982:ASF

[IEE82] IEEE, editor. 23rd annual Symposium on Foundations of Computer Science, November 3–5, 1982, Chicago, Illinois. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1982. CODEN ASF-PDV. ISBN ???? ISSN 0272-5428. LCCN QA76.6 .S95 1982. IEEE catalog no. 82CH1806-9. IEEE Computer Society order no. 440.

IEEE:1983:ASF

[IEE83] IEEE, editor. 24th Annual Symposium on Foundations of Computer Science: November 7–9, 1983, Tucson, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1983. CODEN ASFPDV. ISBN 0-8186-0508-1. ISSN 0272-5428. LCCN QA76.6 .S95 1983. IEEE catalog no. 83CH1938-0.

IEEE:1984:ASF

[IEE84] IEEE, editor. 25th annual Symposium on Foundations of Computer Science, October 24–26, 1984, Singer Island, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

IEEE:1985:ASF

[IEE85] IEEE, editor. 26th annual Symposium on Foundations of Computer Science, October 21–23, 1985, Portland, Oregon. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. CO-DEN ASFPDV. ISBN 0-8186-4644-6 (microfiche), 0-8186-8644-8 (casebound), 0-8186-0644-4 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1985. IEEE catalog no. 85CH2224-4. IEEE Computer Society order no. 644.

IEEE:1986:ASF

[IEE86] IEEE, editor. 27th annual Symposium on Foundations of Computer Science, October 27–29, 1986, Toronto, ON, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. CODEN ASFPDV. ISBN 0-8186-0740-8 (paperback), 0-8186-4740-X (microfiche), 0-8186-8740-1 (casebound). ISSN 0272-5428.

LCCN QA 76 S979 1986. IEEE catalog no. 86CH2354-9. IEEE Computer Society order no. 740.

IEEE:1987:ASF

[IEE87] IEEE, editor. 28th annual Symposium on Foundations of Computer Science, October 12–14, 1987, Los Angeles, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. CODEN ASFPDV. ISBN 0-8186-0807-2, 0-8186-4807-4 (microfiche), 0-8186-8807-6 (casebound). ISSN 0272-5428. LCCN QA 76 S979 1987. IEEE Catalog no. 87CH2471-1. Computer Society order number 807.

IEEE:1988:ASF

[IEE88] IEEE, editor. 29th annual Symposium on Foundations of Computer Science, October 24–26, 1988, White Plains, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

IEEE:1989:ASF

[IEE89] IEEE, editor. 30th annual Symposium on Foundations of Computer Science, October 30–November 1, 1989, Research Triangle Park, North Carolina. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Impagliazzo:1989:OWF

[IL89a] R. Impagliazzo and M. Luby. One-way functions are essential for complexity based cryptography. In IEEE [IEE89], pages 230– 235. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Itkis:1989:PFV

[IL89b] G. Itkis and L. A. Levin. Power of fast VLSI models is insensitive to wires’ thinness. In IEEE [IEE89], pages 402–407. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Impagliazzo:1989:ECS

[IN89] R. Impagliazzo and M. Naor. Efficient cryptographic schemes provably as secure as subset sum. In IEEE [IEE89], pages 236–241. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Ibarra:1984:DSA

[IPK84] O. H. Ibarra, M. A. Palis, and S. M. Kim. Designing systolic algorithms using sequential machines. In IEEE [IEE84], pages 46–55. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Itai:1984:MTA

[IR84] A. Itai and M. Rodeh. The multi-tree approach to reliability in distributed networks. In IEEE [IEE84], pages 137–147. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Impagliazzo:1989:DVS

[IT89] R. Impagliazzo and G. Tardos.

Decision versus search problems in super-polynomial time. In IEEE [IEE89], pages 222–227. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Impagliazzo:1989:HRR

[IZ89] R. Impagliazzo and D. Zuckerman. How to recycle random bits. In IEEE [IEE89], pages 248– 253. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Jacobson:1989:SES

[Jac89] G. Jacobson. Space-efficient static trees and graphs. In IEEE [IEE89], pages 549–554. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Jiang:1989:SNN

[Jia89] T. Jiang. The synchronization of nonuniform networks of finite automata. In IEEE [IEE89], pages 376–381. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Just:1989:GCF

[Jus89] B. Just. Generalizing the continued fraction algorithm to arbitrary dimensions. In IEEE [IEE89], pages 320–324. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Kaminski:1984:MPR

[Kam84] M. Kaminski. Multiplication of polynomials over the ring of integers. In IEEE [IEE84], pages 251–254. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Kharitonov:1989:LBP

[KGY89] M. Kharitonov, A. V. Goldberg, and M. Yung. Lower bounds for pseudorandom number generators. In IEEE [IEE89], pages 242– 247. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Kilian:1988:ZKL

[Kil88] J. Kilian. Zero-knowledge with log-space verifiers. In IEEE [IEE88], pages 25–35. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Kahn:1988:IVB

[KKL88] J. Kahn, G. Kalai, and N. Linial. The influence of variables on Boolean functions. In IEEE [IEE88], pages 68–80. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Klein:1988:EPA

[Kle88] P. N. Klein. Efficient parallel algorithms for chordal graphs. In IEEE [IEE88], pages 150–161. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Kilian:1989:MRZ

[KMO89] J. Kilian, S. Micali, and R. Ostrovsky. Minimum resource zero knowledge proofs. In IEEE [IEE89], pages 474–479. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Kannan:1984:SPA

[KMR84] R. Kannan, G. Miller, and L. Rudolph. Sublinear parallel algorithm for computing the greatest common divisor of two integers. In IEEE [IEE84], pages 7–11. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Kornhauser:1984:CPM

[KMS84] D. Kornhauser, G. Miller, and P. Spirakis. Coordinating pebble motion on graphs, the diameter of permutation groups, and applications. In IEEE [IEE84], pages 241–250. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Khuller:1989:PEP

[KMV89] S. Khuller, S. G. Mitchell, and V. V. Vazirani. Processor efficient parallel algorithms for the two disjoint paths problem, and for finding a Kuratowski homeomorph.

In IEEE [IEE89], pages 300–305. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Ko:1989:CCR

[Ko89] K.-I. Ko. Computational complexity of roots of real functions. In IEEE [IEE89], pages 204–209. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Kock:1988:ISN

[Koc88] R. R. Kock. Increasing the size of a network by a constant factor can increase performance by more than a constant factor. In IEEE [IEE88], pages 221– 230. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Komlos:1984:LVS

[Kom84] J. Komlos. Linear verification for spanning trees. In IEEE [IEE84], pages 201–206. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Kosaraju:1989:ETP

[Kos89a] S. R. Kosaraju. Efficient tree pattern matching. In IEEE [IEE89], pages 178–183. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Kosaraju:1989:PCT

[Kos89b] S. R. Kosaraju. Pipelining computations in a tree of processors. In IEEE [IEE89], pages 184– 189. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Komlos:1988:ECA

[KP88] J. Komlos and R. Paturi. Effect of connectivity in associative memory models. In IEEE [IEE88], pages 138–147. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988.

IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Krentel:1989:SLO

[Kre89] M. W. Krentel. Structure in locally optimal solutions. In IEEE [IEE89], pages 216–221. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Khuller:1989:EPA

[KS89] S. Khuller and B. Schieber. Efficient parallel algorithms for testing connectivity and finding disjoint s–t paths in graphs. In IEEE [IEE89], pages 288–293. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Kirousis:1989:PCS

[KSS89] L. Kirousis, M. Serna, and P. Spirakis. The parallel complexity of the subgraph connectivity problem. In IEEE [IEE89], pages 294– 299. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Kaltofen:1988:CPG

[KT88] E. Kaltofen and B. Trager. Computing with polynomials given by black boxes for their evaluations: greatest common divisors, factorization, separation of numerators and denominators. In IEEE [IEE88], pages 296–305. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Kushilevitz:1989:PCC

[Kus89] E. Kushilevitz. Privacy and communication complexity. In IEEE [IEE89], pages 416–421. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Kuchem:1989:AOT

[KWW89] R. Kuchem, D. Wagner, and F. Wagner. Area-optimal threelayer channel routing. In IEEE [IEE89], pages 506–511. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Landau:1989:SNR

[Lan89] S. Landau. Simplification of nested radicals. In IEEE [IEE89], pages 314–319. CODEN ASF-PDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Levin:1988:HMP

[Lev88] L. A. Levin. Homogeneous measures and polynomial time invariants. In IEEE [IEE88], pages 36– 41. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Liskiewicz:1989:RCA

[LL89] M. Liskiewicz and K. Lorys. On reversal complexity for alternating Turing machines. In IEEE [IEE89], pages 618–623. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Leighton:1989:EMP

[LM89] T. Leighton and B. Maggs. Expanders might be practical: fast algorithms for routing around faults on multibutterflies. In IEEE [IEE89], pages 384–389. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Linial:1989:CDC

[LMN89] N. Linial, Y. Mansour, and N. Nisan. Constant depth circuits, Fourier transform, and learnability. In IEEE [IEE89], pages 574– 579. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Leighton:1988:UPR

[LMR88a] T. Leighton, B. Maggs, and S. Rao. Universal packet routing algorithms. In IEEE [IEE88], pages 256–269. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Linial:1988:RLV

[LMR88b] N. Linial, Y. Mansour, and R. L. Rivest. Results on learnability and the Vapnik–Chervonenkis dimension. In IEEE [IEE88], pages 120– 129. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Leighton:1988:AMF

[LR88] T. Leighton and S. Rao. An approximate max-flow min-cut theorem for uniform multicommodity flow problems with applications to approximation algorithms. In IEEE [IEE88], pages 422–431. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Lovasz:1988:LMF

[LS88] L. Lovasz and M. Saks. Lattices, Möbius functions and communications complexity. In IEEE [IEE88], pages 81–90. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no.

88CH2652-6. Computer Society order no. 877.

Luby:1988:RRP

[Lub88] M. Luby. Removing randomness in parallel computation without a processor penalty. In IEEE [IEE88], pages 162–173. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Li:1989:TLS

[LV89a] M. Li and P. M. B. Vitanyi. A theory of learning simple concepts under simple distributions and average case complexity for the universal distribution. In IEEE [IEE89], pages 34–39. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Linial:1989:GPC

[LV89b] N. Linial and U. Vazirani. Graph products and chromatic numbers. In IEEE [IEE89], pages 124–128. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory.

IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Littlestone:1989:WMA

[LW89] N. Littlestone and M. K. Warmuth. The weighted majority algorithm. In IEEE [IEE89], pages 256–261. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Malitz:1988:GGP

[Mal88] S. M. Malitz. Genus g graphs have pagenumber O(\sqrt{g}). In IEEE [IEE88], pages 458–468. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Mihail:1989:CCM

[Mih89] M. Mihail. Conductance and convergence of Markov chains — a combinatorial treatment of expanders. In IEEE [IEE89], pages 526–531. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no.

89CH2808-4. Computer Society order no. 1982.

Milenkovic:1989:DPG

[Mil89] V. Milenkovic. Double precision geometry: a general technique for calculating line and segment intersections using rounded arithmetic. In IEEE [IEE89], pages 500–505. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Mishra:1984:EAF

[Mis84] B. Mishra. An efficient algorithm to find all ‘bidirectional’ edges of an undirected graph. In IEEE [IEE84], pages 207–216. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Mitchell:1984:SMS

[Mit84] J. C. Mitchell. Semantic models for second-order lambda calculus. In IEEE [IEE84], pages 289–299. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Miller:1989:FPG

[MN89] G. L. Miller and J. Naor. Flow in planar graphs with multiple sources and sinks. In IEEE [IEE89], pages 112–117. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Motwani:1989:PMY

[MNN89] R. Motwani, J. Naor, and M. Naor. The probabilistic method yields deterministic parallel algorithms. In IEEE [IEE89], pages 8–13. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Mehlhorn:1989:CGR

[MNR89] K. Mehlhorn, S. Naher, and M. Rauch. On the complexity of a game related to the dictionary problem. In IEEE [IEE89], pages 546–548. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

McAllester:1988:NFS

[MPS88] D. McAllester, P. Panangaden, and V. Shanbhogue. Nonexpressibility of fairness and signaling. In IEEE [IEE88], pages 377–386. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

MeyeraufderHeide:1984:LSP

[MR84] F. Meyer auf der Heide and R. Reischuk. On the limits to speed up parallel machines by large hardware and unbounded communication. In IEEE [IEE84], pages 56–64. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Motwani:1988:CRG

[MRS88] R. Motwani, A. Raghunathan, and H. Saran. Constructive results from graph minors: linkless embeddings. In IEEE [IEE88], pages 398–409. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Marcotte:1989:FMA

[MS89] O. Marcotte and S. Suri. Fast matching algorithms for points on a polygon. In IEEE [IEE89], pages 60–65. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Moran:1984:ART

[MSM84] S. Moran, M. Snir, and U. Manber. Applications of Ramsey’s theorem to decision trees complexity. In IEEE [IEE84], pages 332–337. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Mansour:1988:LBI

[MST88] Y. Mansour, B. Schieber, and P. Tiwari. Lower bounds for integer greatest common divisor computations. In IEEE [IEE88], pages 54–63. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Mansour:1989:CAS

[MST89] Y. Mansour, B. Schieber, and P. Tiwari. The complexity of approximating the square root. In IEEE [IEE89], pages 325–330. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Maass:1989:CLC

[MT89] W. Maass and G. Turan. On the complexity of learning from counterexamples. In IEEE [IEE89], pages 262–267. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Mulmuley:1984:SCF

[Mul84] K. Mulmuley. A semantic characterization of full abstraction for typed lambda calculi. In IEEE [IEE84], pages 279–288. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Mulmuley:1988:FPP

[Mul88] K. Mulmuley. A fast planar partition algorithm. I. In IEEE [IEE88], pages 580–589. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no.

88CH2652-6. Computer Society order no. 877.

Mulmuley:1989:ORF

[Mul89] K. Mulmuley. On obstructions in relation to a fixed viewpoint. In IEEE [IEE89], pages 592–597. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Munro:1984:IDS

[Mun84] J. I. Munro. An implicit data structure for the dictionary problem that runs in polylog time. In IEEE [IEE84], pages 369–374. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Moses:1988:CTR

[MW88] Y. Moses and O. Waarts. Coordinated traversal: (t + 1)-round Byzantine agreement in polynomial time. In IEEE [IEE88], pages 246–255. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Ng:1989:LBS

[Ng89] C. Ng. Lower bounds for the stable marriage problem and its variants. In IEEE [IEE89], pages 129– 133. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Nisan:1988:HVR

[NW88] N. Nisan and A. Wigderson. Hardness vs. randomness. In IEEE [IEE88], pages 2–11. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Ong:1988:FAM

[Ong88] C.-H. L. Ong. Fully abstract models of the lazy lambda calculus. In IEEE [IEE88], pages 368– 376. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Overmars:1989:OSH

[OS89] M. Overmars and M. Sharir.

Output-sensitive hidden surface removal. In IEEE [IEE89], pages 598–603. CODEN ASFPDV.

ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Overmars:1988:NUB

[OY88] M. H. Overmars and Chee-Keng Yap. New upper bounds in Klee’s measure problem. In IEEE [IEE88], pages 550–556. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Pippenger:1984:PCL

[Pip84] N. Pippenger. Parallel communication with limited buffers. In IEEE [IEE84], pages 127–136. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Pitt:1984:CPI

[Pit84] L. Pitt. A characterization of probabilistic inference. In IEEE [IEE84], pages 485–494. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Plaxton:1989:NCS

[Pla89] C. G. Plaxton. On the network complexity of selection. In IEEE [IEE89], pages 396–401. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Paturi:1984:PCC

[PS84] R. Paturi and J. Simon. Probabilistic communication complexity. In IEEE [IEE84], pages 118–126. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Pacholski:1989:LFC

[PS89] L. Pacholski and W. Szwast. The 0–1 law fails for the class of existential second order Gödel sentences with equality. In IEEE [IEE89], pages 160–163. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Pach:1989:UBN

[PSS89] J. Pach, W. Steiger, and E. Szemeredi. An upper bound on the number of planar k-sets. In IEEE [IEE89], pages 72–79. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Preparata:1988:FDT

[PT88] F. P. Preparata and R. Tamassia. Fully dynamic techniques for point location and transitive closure in planar structures. In IEEE [IEE88], pages 558–567. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Papadimitriou:1984:CTT

[PU84] C. H. Papadimitriou and J. D. Ullman. A communication-time tradeoff. In IEEE [IEE84], pages 84–88. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Renegar:1988:FPA

[Ren88] J. Renegar. A faster PSPACE algorithm for deciding the existential theory of the reals. In IEEE [IEE88], pages 291–295.

CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Ronyai:1989:GGF

[Ron89] L. Ronyai. Galois groups and factoring polynomials over finite fields. In IEEE [IEE89], pages 99– 104. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Ramachandran:1989:OPA

[RR89] V. Ramachandran and J. Reif.

An optimal parallel algorithm for graph planarity. In IEEE [IEE89], pages 282–287. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Russell:1989:FAN

[Rus89] J. R. Russell. Full abstraction for nondeterministic dataflow networks. In IEEE [IEE89], pages 170–175. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Raz:1989:PCC

[RW89] R. Raz and A. Wigderson. Probabilistic communication complexity of Boolean relations. In IEEE [IEE89], pages 562–567. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Safra:1988:CA

[Saf88] S. Safra. On the complexity of ω-automata. In IEEE [IEE88], pages 319–327. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Schapire:1989:SWL

[Sch89] R. E. Schapire. The strength of weak learnability. In IEEE [IEE89], pages 28–33. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Stallmann:1984:APA

[SG84] M. Stallmann and H. N. Gabow.

An augmenting path algorithm for the parity problem on linear matroids. In IEEE [IEE84], pages 217–228. CODEN ASF-PDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Shor:1984:ACA

[Sho84] P. W. Shor. The average-case analysis of some on-line algorithms for bin packing. In IEEE [IEE84], pages 193–200. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Shoup:1988:NAF

[Sho88]  V. Shoup.  New algorithms

for finding irreducible polynomials over finite fields. In IEEE [IEE88], pages 283–290. CODEN ASF-PDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Siegel:1989:UCF

[Sie89] A. Siegel. On universal classes of fast high performance hash functions, their time-space tradeoff, and their applications. In IEEE [IEE89], pages 20–25. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Stern:1987:SLC

[Ste87] J. Stern. Secret linear congruential generators are not cryptographically secure. In IEEE [IEE87], pages 421–426. CODEN ASF-PDV. ISBN 0-8186-0807-2, 0-8186-4807-4 (microfiche), 0-8186-8807-6 (casebound). ISSN 0272-5428. LCCN QA 76 S979 1987. IEEE Catalog no. 87CH2471-1. Computer Society order number 807.

Sullivan:1984:PTA

[Sul84] G. Sullivan. A polynomial time algorithm for fault diagnosability. In IEEE [IEE84], pages 148–156. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Sundar:1989:TTC

[Sun89] R. Sundar. Twists, turns, cascades, deque conjecture, and scanning theorem. In IEEE [IEE89], pages 555–559. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Santha:1984:GQR

[SV84a] M. Santha and U. V. Vazirani. Generating quasi-random sequences from slightly-random sources. In IEEE [IEE84], pages 434–440. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Sedgewick:1984:SPE

[SV84b] R. Sedgewick and J. S. Vitter. Shortest paths in Euclidean graphs. In IEEE [IEE84], pages 417–424. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Tiwari:1984:LBC

[Tiw84] P. Tiwari. Lower bounds on communication complexity in distributed computer networks. In IEEE [IEE84], pages 109–117. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Toda:1989:CPP

[Tod89] S. Toda. On the computational power of PP and (+)P. In IEEE [IEE89], pages 514–519. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Tseng:1989:ELP

[Tse89] W.-G. Tseng. The equivalence and learning of probabilistic automata. In IEEE [IEE89], pages 268–273. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Tarjan:1984:FBC

[TV84] R. E. Tarjan and U. Vishkin. Finding biconnected components and computing tree functions in logarithmic parallel time. In IEEE [IEE84], pages 12–20. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Upfal:1984:HSM

[UW84] E. Upfal and A. Wigderson. How to share memory in a distributed system. In IEEE [IEE84], pages 171–180. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Vaidya:1984:FAA

[Vai84] P. M. Vaidya. A fast approximation algorithm for minimum spanning trees in K-dimensional space. In IEEE [IEE84], pages 403–407. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Vaidya:1989:NAM

[Vai89a] P. M. Vaidya. A new algorithm for minimizing convex functions over convex sets. In IEEE [IEE89], pages 338–343. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Vaidya:1989:SLP

[Vai89b] P. M. Vaidya. Speeding-up linear programming using fast matrix multiplication. In IEEE [IEE89], pages 332–337. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Vazirani:1984:ESP

[VV84] U. V. Vazirani and V. V. Vazirani. Efficient and secure pseudorandom number generation. In IEEE [IEE84], pages 458–463. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

vonzurGathen:1984:PP

[vzG84] J. von zur Gathen. Parallel powering. In IEEE [IEE84], pages 31–36. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

vonzurGathen:1989:TPP

[vzG89] J. von zur Gathen. Testing permutation polynomials. In IEEE [IEE89], pages 88–92. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Yao:1988:NOT

[Yao88] A. C.-C. Yao. Near-optimal timespace tradeoff for element distinctness. In IEEE [IEE88], pages 91– 97. CODEN ASFPDV. ISBN 0-8186-0877-3 (paperback), 0-8186-4877-5 (microfiche), 0-8186-8877-7 (hard). ISSN 0272-5428. LCCN QA 76 S979 1988. IEEE catalog no. 88CH2652-6. Computer Society order no. 877.

Yao:1989:LBA

[Yao89] A. C.-C. Yao. Lower bounds for algebraic computation trees with integer inputs. In IEEE [IEE89], pages 308–313. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.